2x-3y=-11
5x+2y=1

Answer by Alan3354(69443)

2x-3y=-11
5x+2y=1
----------
|2 -3 -11|
|5 +2 +1|
-------------
Det = 4 + 15 = 19
--------
-x*Det = -3 + 22 = 19
x = -1
-------
y*Det = 2 + 55 = 57
y = 3